OpenStudy (imamermaid29):

What is the solution to the system of equations? x + y = 2 x - y = 4 ( , )

OpenStudy (anonymous):

x is 3, y is -1 3 + (-1) = 3-1 = 2 3 - (-1) = 3+1 = 4

OpenStudy (anonymous):

you really just have to guess and check

OpenStudy (anonymous):

imsorry, i don't know how to explain this to you...

OpenStudy (anonymous):

lol

OpenStudy (imamermaid29):

thank you :) @sarenemun

OpenStudy (anonymous):

:) no problem i just wished i knew how i could explain how i got the answer...